# ContextGPT: Infusing LLMs Knowledge into Neuro-Symbolic Activity Recognition Models

### v1

@@ 3.3.2. Context2Text
An example of application of Context2Text is depicted in Figure 5.
Figure 5. Context2Text applied to a specific context situation.
3.3.3. Example pool
As previously mentioned, LLMs benefit from including a few examples in the prompt, showing the output that should be associated with specific inputs. In our case, an example represents a context situation and the corresponding consistent activities. We assume that the Prompt Engineer is sufficiently knowledgeable in the activity domain and, given a context description, they are capable of determining which activities are consistent. For each activity $a$, the Prompt Engineer populates the pool of examples $P$ with the following strategy:
  • Define a set of examples $E_{a}$ (i.e. a set of tuples ¡$context$, $consistentActivities$¿) referring to context situations that are uncommon but possible for the activity $a$ (the context may be common or uncommon for the other consistent activities)
@@ 3.3.3. Example pool
3.3.4. Example Selection
Since the number of examples in the pool is not fixed, including all of them in the prompt may result in increased costs, especially when using external services whose price depends on the prompt’s length. Moreover, there are often limits to the LLM’s input size. Hence, if the number of examples is high, they may exceed the model’s capacity. ContextGPT employs a method to include in the prompt only those examples that are similar to the input context $C^{w}$, and its pseudo-code is outlined in Algorithm 1.
$S\leftarrow\emptyset$ ;
$C^{w}\leftarrow$ current user’s context;
$T^{C}\leftarrow Context2Text(C^{w})$ ;
$E^{C}\leftarrow$ compute embeddings of $T^{C}$ using a LLM ;
forall $e\in P$ do
       $T^{e}\leftarrow Context2Text(e.context)$;
       $E^{e}\leftarrow$ compute embeddings of $T^{e}$ using a LLM ;
       $s_{e}\leftarrow\frac{E^{C}\cdot E^{e}}{|E^{C}||E^{e}|}$ ;
       if $s_{e}>k$ then
             $S\leftarrow S\bigcup\{e\}$;
       end if
end forall
return $S$
Algorithm 1 Example Selection
Specifically, we use a pre-trained LLM to extract text embeddings from $C^{w}$ and all the examples in the pool using their description in natural language obtained with Context2Text. Then, we use cosine similarity to compute the similarity of each example with $C^{w}$, and we include in the prompt only the examples with a similarity higher than a threshold $k$. This threshold is determined empirically, and we will show its impact on the recognition rate in Section 4.
3.4. Post-processing
@@
Figure 6. An example of LLM output
As required by the system message (Section 3.3.1), besides explaining the reasoning process, the output includes the list of consistent activities in square brackets. Hence, it is possible to easily extract this list $L$ using regular expressions and transform it into a binary vector $b=[b_{1},b_{2},\dots,b_{n}]$ where $b_{i}$ is $1$ if the activity $a_{i}\in L$ (i.e., the activity $a_{i}$ is consistent with $C^{w}$ according to the LLM), 00 otherwise. In the following, we will refer to $b$ as the consistency vector. Note that in the actual implementation of ContextGPT there is a cache mechanism that avoids recomputing the consistency vector when $C^{w}$ has already been processed. Consistency vectors are used to infuse knowledge inside the NeSy model.
3.5. Infusing knowledge into Neuro-Symbolic model
While ContextGPT is agnostic with respect to the Neuro-Symbolic approach for Context-Aware HAR, in this work we use a specific knowledge infusion approach proposed in the general AI community (Sheth et al., 2019; Kursuncu et al., 2019). This method relies on a hidden layer in the Deep Neural Network (DNN) in charge of infusing knowledge in the latent space. An adaptation for Context-Aware HAR named NIMBUS has been recently proposed in the literature, exhibiting promising recognition rates (Arrotta et al., 2022). In NIMBUS, knowledge is infused from a manually defined ontology of activity and contexts. In this work, we adapted NIMBUS to fetch knowledge from ContextGPT.
@@ 3.5. Infusing knowledge into Neuro-Symbolic model
Figure 7. Infusing ContextGPT into the Symbolic Features approach
4. Experimental Evaluation
4.1. Datasets
In the following, we describe the two publicly available datasets we used to evaluate ContextGPT. Both datasets contain real data with the former obtained in a realistic but scripted setting, and the latter collected in a in-the-wild setting. They include two types of data collected from mobile devices: a) raw inertial sensor data and b) pre-processed high-level context data.
4.1.1. DOMINO
@@
The DOMINO (Arrotta et al., 2023b) dataset includes data from $25$ subjects wearing a smartwatch on the wrist of their dominant hand and a smartphone in their pants front pocket. Both devices gathered raw sensor data from inertial sensors (accelerometer, gyroscope, and magnetometer) and a wide variety of high-level context data. This dataset includes nearly $9$ hours of labeled data (approximately $350$ activity instances) covering $14$ distinct activities: walking, running, standing, lying, sitting, stairs up, stairs down, elevator up, elevator down, cycling, moving by car, sitting on transport, standing on transport, and brushing teeth. The DOMINO dataset was collected in a scripted setting: the volunteers were instructed to perform sequences of indoor/outdoor activities, even though without specific guidance on their execution.
DOMINO included the following pre-processed high-level context information:
  • User’s height variations (discretized: negative, null, positive)
  • User’s speed variations (discretized: null, low, medium, and high)
  • Whether the user is indoors or outdoors
  • Semantic locations (Home, Office, University, Mall, Station, Museum, Gym, Shop, Bar, Restaurant, Barbershop, Bank, Church)
  • Weather condition (Sunny, Rainy, Misty, Cloudy, Drizzly, Stormy)
  • Whether the user is following a public transportation route
Overall, the DOMINO dataset includes data collected in $412$ unique context conditions.
4.1.2. ExtraSensory
ExtraSensory (Vaizman et al., 2018) was collected in an unscripted in-the-wild setting from $60$ subjects. Similarly to DOMINO, users wore a smartwatch on the wrist of their dominant hand and a smartphone in their pants front pocket. ExtraSensory includes approximately $300$,$000$ minutes of labeled data, including $51$ distinct labels self-reported by the users. These labels encode both high-level context information (e.g., at home, with friends, phone in bag, phone is charging) and performed activities (e.g., sitting, bicycling). Considering inertial sensor data, each smartphone collected raw data from the accelerometer, gyroscope, and magnetometer; on the other hand, the smartwatch only collected raw data from the accelerometer.
Since it was collected in the wild, ExtraSensory is widely considered a challenging benchmark. Existing works on this dataset report low recognition rates although considering small subsets of activities (Arrotta et al., 2023a; Tarafdar and Bose, 2021; Cruciani et al., 2020).
In this paper, we pre-process the dataset consistently with previous works in context-aware HAR (Arrotta et al., 2023a). We consider the following $7$ activities: bicycling, lying down, moving by car, on transport, sitting, standing, and walking. All the context information in the dataset that could be easily collected by mobile devices is provided as input to the Neuro-Symbolic model (e.g., audio level, screen brightness, ringer mode, etc.).
However, based on preliminary experiments, we provide to the LLM only the context information where common-sense knowledge can be practically used to reason about the target physical activities:
  • Whether the user is indoors or outdoors
  • Semantic location (Home, Workplace, School, Gym, Restaurant, Shopping, Bar, Beach)
  • User’s speed (null, low, medium, and high)
  • User’s movements diameter (null, low, medium, and high)
  • Whether the user is following a public transportation route
Overall, the ExtraSensory data presents $144$ unique context conditions for the LLM. This number is significantly lower compared to DOMINO, due to a reduced number of context variables and target activities.
4.2. Experimental setup
@@
We implemented a working prototype of ContextGPT as well as the Neuro-Symbolic model explained in Section 3.5 using the Python language (version 3.12). We run our experiments on a machine of our department running Ubuntu 20.04.4 LTS and equipped with an AMD EPYC Processor x86-64, an NVIDIA A100 GPU (80 GB), with $43.5$ GBs RAM allocated. In the following, we provide details about our experimental setup.
4.2.1. Large Language Models
@@
The pre-trained LLM we used in our experiments is gpt-3.5-turbo by OpenAI, queried through its API through the Python OpenAI package (version 0.28.1). We set the temperature of the model to 00 to ensure a mostly deterministic output since our task does not require leveraging the model’s creativity. The pre-trained LLM model we adopted to compute example embeddings and thus computing similarity (see Section 3.3.4) is all-MiniLM-L6-v2111https://huggingface.co/sentence-transformers/all-MiniLM-L6-v2.
4.2.2. Generating examples
A member of our research group, without experience with the specific datasets used in this work, created the examples following the strategy described in Section 3.3.4. To simplify the example creation task and to automatically generate examples in natural language, we developed a simple examples creation tool (see Figure 8).
Figure 8. A screenshot of our tool to create examples
@@ 4.2.2. Generating examples
Since each dataset has different sets of activities and context variables, examples were created separately for DOMINO and ExtraSensory. The output of this process is $21$ examples for DOMINO and $12$ for ExtraSensory.
4.2.3. Model and Hyper-parameters
For the NeSy model, we use the implementation of NIMBUS proposed in (Arrotta et al., 2023a) [^2] Note that in (Arrotta et al., 2023a) NIMBUS is named Symbolic Features.. Smartphone inertial sensor data are provided to a sequence of three convolutional layers, each consisting of $32$, $64$, and $96$ filters with corresponding kernel sizes of $24$, $16$, and $8$. These layers are interleaved with max-pooling layers with a pool size of $4$. Then, the flow continues with a global max pooling layer and a fully connected layer with $128$ neurons. In parallel on another channel, the smartwatch inertial sensor data is processed by an identical sequence of layers with a small difference: the kernel size of the first three convolutional layers is $16$, $8$, and $4$, respectively. In parallel, high-level context data (represented with one-hot encoding) is provided to a single fully connected layer with $8$ neurons. The features extracted from these three parallel streams are merged with the consistency vector from ContextGPT using a concatenation layer (i.e., the knowledge-infusion layer presented in Section 3.5). Then, a dropout layer (with a rate of $0.1$) is introduced for regularization, followed by a fully connected layer of $256$ neurons to extract correlations among the features from the different data sources and the consistency vector. Finally, a softmax layer returns the probability distribution over the possible activities.
We used the same hyper-parameters proposed in (Arrotta et al., 2023a):
  • Segmentation window: $4$ seconds
  • Optimizer: Adam
  • Loss: Categorical cross-entropy
  • Learning rate: $0.001$
  • Batch size: $32$
  • Epochs: $200$
  • Early stopping patience on the validation loss: $5$
4.3. Evaluation methodology
4.3.1. Baselines
@@
  • Ontology: the NIMBUS approach originally proposed in (Arrotta et al., 2022), where knowledge is infused from an ontology encoding relationships between activities and contexts. Hence, we use the ontology adopted in that work for the DOMINO dataset and its adaptation to the ExtraSensory dataset recently proposed in (Arrotta et al., 2023a).
4.3.2. Evaluation Strategy
We evaluated the recognition rate of the HAR models by using a leave-$n$-users-out cross-validation technique. At each fold, $n$ users were designated for the test set, while the remaining users were split between the training ($90\%$) and validation ($10\%$) sets. Specifically, for the DOMINO dataset we set $n=1$, while for the Extrasensory dataset, consistently with other works in the literature (Cruciani et al., 2020), we set $n=5$.
We simulated several data scarcity scenarios by downsampling the available training data at each fold. At each iteration, we used the test set to evaluate the recognition rate in terms of macro F1 score. To ensure robust results, we conducted each experiment $5$ times, calculating the average F1 score.
4.4. Results
4.4.1. Comparison with the baselines
Figures 9 and  10 show, for both datasets, the impact of infusing ContextGPT knowledge (using the $k$ value leading to the best results) compared to the baselines.
Figure 9. DOMINO: Knowledge infusion with ContextGPT compared to the baselines. This plot shows the F1 score in different data scarcity scenarios.
Figure 10. ExtraSensory: ContextGPT compared to the baselines. This plot shows the F1 score in different data scarcity scenarios.
@@ 4.4.1. Comparison with the baselines
Consistently with existing works (Cruciani et al., 2020; Tarafdar and Bose, 2021; Arrotta et al., 2023a), the in-the-wild nature of ExtraSensory leads to relatively low recognition rates. Moreover, on this dataset, increasing the percentage of training data slightly degrades the recognition rate of both knowledge infusion approaches. The reason may be due to the underlying NeSy model, in which raw sensor data may overshadow the infused knowledge when the training set is large. Since DOMINO is significantly smaller, we do not observe the same phenomenon on this dataset.
@@
4.4.2. Impact of examples selection
Figures 11 and 12 illustrate the distribution of the number of selected examples by varying $k$ for DOMINO and ExtraSensory, respectively.
Figure 11. DOMINO: Distribution of the number of selected examples by varying $k$.
Figure 12. ExtraSensory: Distribution of the number of selected examples by varying $k$.
We observe that in both cases $k=0.75$ is the lowest threshold value significantly restricting the number of selected examples, while lower values lead to a higher number of examples.
In Figure 9, the best results on DOMINO were achieved with low values of $k$ (i.e., ranging from 00 to $0.25$) and thus a high number of examples. This is due to the fact that, in this dataset, there is a high number of possible context conditions. Thus, the LLM benefits from more examples describing the required task.
@@ 4.4.2. Impact of examples selection
On the other hand, in Figure 10, the best results on the ExtraSensory dataset are associated with higher values of $k$ (i.e., ranging from $0.25$ to $0.95$) and thus selecting a smaller number of examples. In this case, since the number of possible context conditions is significantly lower compared to DOMINO, a few examples (similar to the input context) provide the LLM model a better guidance on the required task.
In the following, we show the impact of example selection on the recognition rate. Figure 13 and 14 show how $k$ impacts the recognition rate on each dataset when only the $10\%$ of the training data is available. We observe that the DOMINO dataset in this case benefits from $k=0$ (i.e., using all the examples), while higher values of $k$ lead to worse recognition rates. Nonetheless, it is interesting to note that even $k=1$ (i.e., no examples) leads to outperforming the No knowledge baseline by ${}\approx+4\%$.
On the other hand, the best recognition rates reached on the ExtraSensory dataset in the same data scarcity scenario are associated with $k=0.5$ and $k=0.75$. Notably, $k=0.75$ leads to a significantly smaller number of examples than $k=0.5$ and to a similar recognition rate.
Our results on these datasets suggest that the higher the number of context variables and activities, the higher the number of examples needed in the prompt to maximize the recognition rate. We believe that this reliance on examples may be mitigated by future more comprehensive LLMs and with fine-tuning.
Figure 13. DOMINO: Impact of k on macro F1 with 10% of the training set
Figure 14. ExtraSensory: Impact of k on macro F1 with 10% of the training set
While examples play a major role in improving the recognition rate for data scarcity scenarios, they do not lead to significant improvement when sufficient amounts of training data are available. This phenomenon is depicted in Figures 15 and 16, where we show the impact of $k$ when the $50\%$ of training data is available.
Figure 15. DOMINO: Impact of k on macro F1 with 50% of the training set
Figure 16. ExtraSensory: Impact of k on macro F1 with 50% of the training set
When enough training data is available, the knowledge gaps in the LLM models addressed by examples are probably learned from training data. Nonetheless, infusing knowledge from ContextGPT still outperforms the No knowledge baseline.
5. LLM vs. Ontology for Neuro-Symbolic HAR
Our results show that by using pre-trained LLMs instead of ontologies in NeSy HAR systems, we can reach similar (and sometimes better) recognition results.
@@
Then we may wonder why LLMs are a more promising approach for future real deployment of these systems, since we already have some ontologies like the ones used in our work, while for LLMs some prompt engineering work is required.
@@
Considering the extension of these systems to a large class of human activities and different context situations, there are clear limitations for ontologies. To the best of our knowledge, we are not aware of publicly available ontologies offering a comprehensive coverage of all possible human activities and context situations. Hence, significant human effort would be required to extend and adapt the ontology to new datasets with different sets of activities and context data.
Indeed, extending an ontology means defining complex knowledge relationships between activities and contexts, it requires skills in the logic-based formalism model (e.g., OWL2 in the case of ontologies) and significant expertise in the HAR domain. This task is also usually assigned to a single knowledge engineer or small team with high risks of incompleteness.
In our case, adapting ContextGPT to a new dataset only requires using natural language to adjust the system message on the target activities, extending the Context2Text module to map new contexts to natural language descriptions, and generating a new pool of examples.
@@
However, a significant disadvantage of LLMs compared to ontologies is the absence of real semantic reasoning, since the output is based on data-driven text generation. Hence, there may be contradictions and/or model hallucinations that we would not experience by using rigorous logic systems. For instance, in one instance where the user was moving slowly, the model considered as possible standing with the following motivation:“Possible, as the user is moving at a relatively slow pace”. While hallucinations may be mitigated by more advanced LLMs models (e.g., GPT-4) we believe that knowledge infusion needs to cope with possibly noisy information.
Finally, we investigate the discrepancies between the sets of consistent activities obtained by ContextGPT with the ones generated by the ontology. Specifically, given a context condition, we compute the set $L$ of activities considered consistent by the LLM in ContextGPT and $O$ as the set of activities considered consistent by the ontology. Hence, we define the following metrics:
  • $L2O$ Inclusion: the proportion of activities considered as consistent by ContextGPT that are also consistent for the ontology, computed as $\frac{|L\bigcap O|}{|L|}$.
  • $O2L$ Inclusion: the proportion of activities considered as consistent by the ontology that are also consistent for ContextGPT, computed as $\frac{|L\bigcap O|}{|O|}$.
Note that the low values of $L2O$ Inclusion imply that ContextGPT includes several activities that are not considered consistent by the ontology. Conversely, low values of $O2L$ Inclusion imply that ContextGPT does not consider as consistent several activities that are deemed consistent by the ontology.
Figure 17. DOMINO: Average $L2O$ and $O2L$ inclusion metrics varying $k$.
Figure 18. ExtraSensory: Average $L2O$ and $O2L$ inclusion metrics varying $k$.
Figure 17 shows, for the DOMINO dataset, the average scores of $L2O$ and $O2L$ inclusion metrics at different values of $k$. By increasing $k$, we observe that the $O2L$ metric is the one associated with the more drastic decrease, thus denoting that considering only a small number of examples leads ContextGPT to be restrictive on the consistent activities compared to the ontology. Moreover, high values of $k$ are also associated with lower values of $L2O$. This decrease in both metrics indicates a significant discrepancy between the activities considered consistent by ContextGPT and those considered consistent by the ontology. We hypothesize this may be correlated with the results depicted in Figure 13, where a substantial number of examples was required by ContextGPT to achieve results comparable to those obtained by Ontology on the DOMINO dataset.
Figure 18 shows the impact of the inclusion metrics on ExtraSensory. In this case, increasing $k$ significantly impacts only the $O2L$ Inclusion metric, with ContextGPT considering a restricted set of consistent activities compared to the ontology. On the other hand, the $L2O$ Inclusion metric exhibits only a slight decrease since the activities consistent for ContextGPT are often consistent also for the ontology. We hypothesize that such high values of the $L2O$ metric are because, on this dataset, reducing the number of examples does not lead to significant LLM’s hallucinations. This seems to be consistent with the results previously depicted in Figure 14.
6. Conclusions and Future Work

### v2

@@ III-C2 Context2Text
III-C3 Example pool
LLMs benefit from including in the prompt a few examples showing the output that should be associated with specific inputs. In our case, an example represents a context situation and the corresponding consistent activities. We assume that the Prompt Engineer is sufficiently knowledgeable in the activity domain and, given a context description, they are capable of determining which activities are consistent. For each activity $a$, the Prompt Engineer populates the pool of examples $P$ with the following strategy:
  • Define a set of examples $E_{a}$ (i.e. a set of tuples $\langle context$, $consistentActivities\rangle$) referring to context situations that are uncommon but possible for the activity $a$ (the context may be common or uncommon for the other consistent activities)
@@ III-C3 Example pool
III-C4 Example Selection
Since the number of examples in the pool is not fixed, including all of them in the prompt may result in increased costs, especially when using external services whose price depends on the prompt’s length. Moreover, there are often limits to the LLM’s input size.
ContextGPT employs a method to include in the prompt only those examples that are similar to the input context $C^{w}$. Specifically, we use a pre-trained LLM to extract text embeddings from $C^{w}$ and all the examples in the pool using their description in natural language obtained with Context2Text. Then, we use cosine similarity to compute the similarity of each example with $C^{w}$, and we include in the prompt only the examples with a similarity higher than a threshold $k$. This threshold is determined empirically, and we will show its impact on the recognition rate in Section IV.
III-D Post-processing
@@
Figure 4: An example of LLM output
As required by the system message (Section III-C1), besides explaining the reasoning process, the output includes the list of consistent activities in square brackets. Hence, it is possible to easily extract this list $L$ using regular expressions and transform it into a binary vector $b=[b_{1},b_{2},\dots,b_{n}]$ where $b_{i}$ is $1$ if the activity $a_{i}\in L$ (i.e., the activity $a_{i}$ is consistent with $C^{w}$ according to the LLM), 00 otherwise. In the following, we will refer to $b$ as the consistency vector. Note that in the actual implementation of ContextGPT there is a cache mechanism that avoids recomputing the consistency vector when $C^{w}$ has already been processed. Consistency vectors are used to infuse knowledge inside the NeSy model.
III-E Infusing knowledge into Neuro-Symbolic model
While ContextGPT is agnostic with respect to the Neuro-Symbolic approach for Context-Aware HAR, in this work we use a specific knowledge infusion approach proposed in the general AI community [29, 30]. This method relies on a hidden layer in the Deep Neural Network (DNN) in charge of infusing knowledge in the latent space. An adaptation for Context-Aware HAR named NIMBUS has been recently proposed in the literature, exhibiting promising recognition rates [6]. In NIMBUS, knowledge is infused from a manually defined ontology of activity and contexts. In this work, we adapted NIMBUS to fetch knowledge from ContextGPT.
@@ III-E Infusing knowledge into Neuro-Symbolic model
Figure 5: Infusing ContextGPT into the Symbolic Features approach
IV Experimental Evaluation
IV-A Datasets
IV-A1 DOMINO
@@
The DOMINO [28] dataset includes data from $25$ subjects wearing a smartwatch on the wrist of their dominant hand and a smartphone in their pants front pocket. Both devices gathered raw sensor data from inertial sensors and a wide variety of high-level context data. This dataset includes nearly $9$ hours of labeled data (approximately $350$ activity instances) covering $14$ distinct activities: walking, running, standing, lying, sitting, stairs up, stairs down, elevator up, elevator down, cycling, moving by car, sitting on transport, standing on transport, and brushing teeth. The DOMINO dataset was collected in a scripted setting: the volunteers were instructed to perform sequences of indoor/outdoor activities, even though without specific guidance on their execution.
DOMINO included several pre-processed high-level context information, like the semantic loation, weather conditoin, whether the user is following a public transportation route, user’s height and speed variations, etcetera. Overall, the DOMINO dataset includes data collected in $412$ unique context conditions.
IV-A2 ExtraSensory
ExtraSensory [31] was collected in an unscripted in-the-wild setting from $60$ subjects. Similarly to DOMINO, users wore a smartwatch on the wrist of their dominant hand and a smartphone in their pants front pocket. ExtraSensory includes approximately $300$,$000$ minutes of labeled data, including $51$ distinct labels self-reported by the users. These labels encode both high-level context information (e.g., at home, with friends, phone in bag, phone is charging) and performed activities (e.g., sitting, bicycling). Since it was collected in the wild, ExtraSensory is widely considered a challenging benchmark. Existing works on this dataset report low recognition rates although considering small subsets of activities [18, 32]. In this paper, we pre-process the dataset consistently with previous works in context-aware HAR [18]. We consider the following $7$ activities: bicycling, lying down, moving by car, on transport, sitting, standing, and walking. All the context information in the dataset that could be easily collected by mobile devices is provided as input to the Neuro-Symbolic model (e.g., audio level, screen brightness, ringer mode, etc.). However, based on preliminary experiments, we provide to the LLM only the context information where common-sense knowledge can be practically used to reason about the target physical activities: the user’s semantic location, their speed, movements diameters, and whether they are following a public transportation route. Overall, the ExtraSensory data presents $144$ unique context conditions for the LLM. This number is significantly lower compared to DOMINO, due to a reduced number of context variables and target activities.
IV-B Experimental setup
@@
We implemented a working prototype of ContextGPT as well as the Neuro-Symbolic model explained in Section III-E using the Python language (version 3.12). We run our experiments on a machine of our department running Ubuntu 20.04.4 LTS and equipped with an AMD EPYC Processor x86-64, an NVIDIA A100 GPU (80 GB), with $43.5$ GBs RAM allocated.
@@
The pre-trained LLM we used in our experiments is gpt-3.5-turbo by OpenAI, queried through its API through the Python OpenAI package (version 0.28.1). We set the temperature of the model to 00 to ensure a mostly deterministic output since our task does not require leveraging the model’s creativity. The pre-trained LLM model we adopted to compute example embeddings and thus computing similarity (see Section III-C4) is all-MiniLM-L6-v2111https://huggingface.co/sentence-transformers/all-MiniLM-L6-v2. Considering the NeSy model, we adopted the NIMBUS implementation proposed in [18] [^2] Note that in [18] NIMBUS is named Symbolic Features. (with the same hyper-parameters).
@@
Finally, a member of our research group (without experience with the specific datasets used in this work) created the examples for few-shot-prompting following the strategy described in Section III-C4. Since each dataset has different sets of activities and context variables, examples were created separately for DOMINO and ExtraSensory. The output of this process is $21$ examples for DOMINO and $12$ for ExtraSensory.
IV-C Evaluation methodology
IV-C1 Baselines
@@
  • Ontology: the NIMBUS approach originally proposed in [6], where knowledge is infused from an ontology encoding relationships between activities and contexts. We use the ontology adopted in that work for the DOMINO dataset and its adaptation to the ExtraSensory dataset recently proposed in [18].
(a) DOMINO
(b) ExtraSensory
Figure 6: Knowledge infusion with ContextGPT compared to the baselines. This plot shows the F1 score in different data scarcity scenarios.
IV-C2 Evaluation Strategy
We evaluated the recognition rate of the HAR models by using a leave-$n$-users-out cross-validation technique. At each fold, $n$ users were designated for the test set, while the remaining users were split between the training ($90\%$) and validation ($10\%$) sets. Specifically, for the DOMINO dataset we set $n=1$, while for the Extrasensory dataset, consistently with other works in the literature [32], we set $n=5$. We simulated several data scarcity scenarios by downsampling the available training data at each fold. At each iteration, we used the test set to evaluate the recognition rate in terms of macro F1 score. To ensure robust results, we conducted each experiment $5$ times, calculating the average F1 score.
IV-D Results
Figures 6(a) and  6(b) show, for both datasets, the impact of infusing ContextGPT knowledge (using the $k$ value leading to the best results) compared to the baselines.
@@ IV-D Results
Consistently with existing works [32, 18], the in-the-wild nature of ExtraSensory leads to relatively low recognition rates. Moreover, on this dataset, increasing the percentage of training data slightly degrades the recognition rate of both knowledge infusion approaches. The reason may be due to the underlying NeSy model, in which raw sensor data may overshadow the infused knowledge when the training set is large. Since DOMINO is significantly smaller, we do not observe the same phenomenon on this dataset.
@@
The best results on DOMINO were achieved with low values of $k$ (i.e., ranging from 00 to $0.25$) and thus a high number of examples. This is due to the fact that, in this dataset, there is a high number of possible context conditions. Thus, the LLM benefits from more examples describing the required task. On the ExtraSensory dataset, the best results are associated with higher values of $k$ (i.e., ranging from $0.25$ to $0.95$) and thus selecting a smaller number of examples. In this case, since the number of possible context conditions is significantly lower compared to DOMINO, a few examples (similar to the input context) provide the LLM model a better guidance on the required task.
V LLM vs. Ontology for Neuro-Symbolic HAR
@@
Our results show that by using pre-trained LLMs instead of ontologies in NeSy HAR systems, we can reach similar (and sometimes better) recognition results. Then we may wonder why LLMs are a more promising approach for future real deployment of these systems, since we already have some ontologies like the ones used in our work, while for LLMs some prompt engineering work is required.
@@
Considering the extension of these systems to a large class of human activities and different context situations, there are clear limitations for ontologies. To the best of our knowledge, we are not aware of publicly available ontologies offering a comprehensive coverage of all possible human activities and context situations. Hence, significant human effort would be required to extend and adapt the ontology to new datasets with different sets of activities and context data. Indeed, extending an ontology means defining complex knowledge relationships between activities and contexts, it requires skills in the logic-based formalism model (e.g., OWL2 in the case of ontologies) and significant expertise in the HAR domain. This task is also usually assigned to a single knowledge engineer or small team with high risks of incompleteness. In our case, adapting ContextGPT to a new dataset only requires using natural language to adjust the system message on the target activities, extending the Context2Text module to map new contexts to natural language descriptions, and generating a new pool of examples.
@@
However, a significant disadvantage of LLMs compared to ontologies is the absence of real semantic reasoning, since the output is based on data-driven text generation. Hence, there may be contradictions and/or model hallucinations that we would not experience by using rigorous logic systems. For instance, in one instance where the user was moving slowly, the model considered as possible standing with the following motivation:“Possible, as the user is moving at a relatively slow pace”. While hallucinations may be reduced by more advanced LLMs models (e.g., GPT-4), they may still occur. One of the advantages of NeSy models is that they can robustly cope with noise in the infused knowledge [18].
knowledge infusion needs to cope with possibly noisy information.
VI Conclusions and Future Work